the perimeter of a rectangle is 82 yards. The width is 15 yards less than the length. Find the lenght and the width of the rectangle.
Answer by 303795(562)

Call the length of the rectangle l. the width is 15 yards less than the length so the width is l - 15.
The perimeter of the rectangle = length + length + width + width
Using the length (l) and width (l - 15) then the perimeter can be found by
Perimeter = l + l + l -15 + l - 15 = 4l - 30
This equals 82 yards so

4l - 30 = 82 Add 30 to each side of the equation
4l = 82+30 = 112
4l = 112 Divide each side of the equation by 4
l = 112 / 4 = 28 Yards

The width is 15 less than this ie 28 - 15 = 13 Yards

(Be careful with the letter l which I have used in the example. It can look like a number one.)